# GeneMamba: An Efficient and Effective Foundation Model on Single Cell Data

### v2

@@ 1 Introduction
Single-cell RNA sequencing (scRNA-seq) has fundamentally transformed our ability to study cellular heterogeneity and dynamics by enabling high-resolution profiling of gene expression at the individual cell level korsunsky2019fast; theodoris2023transfer. This technology allows researchers to dissect complex cellular compositions within tissues, trace differentiation trajectories, and identify rare cell populations that remain undetectable with bulk RNA sequencing shen2023generative; wen2023single; deconinck2021recent. The rich transcriptomic insights provided by scRNA-seq have catalyzed significant advancements in developmental biology, disease modeling, and drug discovery gu2023mamba; yang2024genecompass; li2024single.
@@
Despite these advantages, the computational analysis of scRNA-seq data presents formidable challenges due to its high dimensionality, inherent sparsity, and technical noise du2019gene2vec; zhao2023large. Addressing these challenges requires robust computational models capable of capturing biologically meaningful patterns while efficiently handling the scale and variability of single-cell data. Transformer-based architectures, such as scBERT yang2022scbert and scGPT cui2024scgpt, have emerged as powerful tools in this domain, demonstrating strong performance in tasks such as cell type classification, gene expression imputation, and differential expression analysis bian2024scmulan; wen2023cellplm; szalata2024transformers.
@@
However, transformers exhibit critical limitations when applied to scRNA-seq data. The quadratic complexity of their self-attention mechanism constrains their scalability for long sequences, which are characteristic of single-cell transcriptomes vaswani2017attention; choromanski2020rethinking. Moreover, transformers often struggle to effectively capture long-range dependencies, which are essential for modeling gene regulatory interactions and cell state transitions dao2022flashattention; chen2023transformer. These limitations have driven the exploration of alternative architectures, among which state space models (SSMs) have emerged as promising solutions, offering improved efficiency and scalability for processing long sequences gu2023mamba; mamba2.
@@
In this study, we introduce GeneMamba, a novel state space model designed to efficiently train large-scale cell models on scRNA-seq data. SSM-based models have demonstrated competitive or superior performance compared to transformers while significantly reducing computational overhead shen2022universal; gong2024xtrimogene; ryu2023integration. Building upon this foundation, GeneMamba incorporates bidirectional computation, enhancing its ability to capture both upstream and downstream contextual dependencies, thereby improving its performance in tasks requiring global contextual awareness liu2024bidirectional; liang2024bi. We validate GeneMamba across a diverse set of applications, including multi-batch integration, cell type annotation, and gene-gene correlation analysis. Our experimental results demonstrate substantial improvements in computational efficiency and predictive performance compared to existing models hao2024large; sun2024bidirectional; petegrosso2020machine. Furthermore, to assess its scalability, we conduct a comprehensive performance analysis on various variations of GeneMamba (Appendix F), highlighting its advantages in real-world biological applications.
@@ 2.1 Discretization of Gene Expression
Bin-based discretization, as used by scBERT yang2022scbert, scGPT cui2024scgpt, and scMulan bian2024scmulan, groups expression values into predefined bins. This approach preserves absolute value distributions and simplifies sequence modeling, but may introduce information loss, particularly for genes with subtle but biologically significant expression differences. Additionally, binning can be sensitive to parameter selection, affecting downstream results.
Value projectionszalata2024transformers, adopted by scFoundation hao2024large and its backbone model xTrimoGene gong2024xtrimogene, projects gene expression values into continuous embeddings rather than discrete categories. This method maintains full data resolution by applying a linear transformation to the gene expression vector, which is then combined with gene-specific embeddings. However, the use of continuous embeddings diverges from traditional tokenization strategies in NLP-based transformers, and its impact on model performance remains an open question.
Rank-based discretization, utilized by Geneformer theodoris2023transfer, GeneCompass yang2024genecompass, tGPTshen2023generative and LangCell zhao2024langcell, transforms gene expression values into ordinal rankings. This approach effectively captures relative expression levels and is more robust to batch effects and noise. Our method is based on rank-based discretization, as employed in Geneformer, which aligns naturally with biological processes such as regulatory interactions.
@@ 2.2 Model Architectures for Single-Cell Analysis
Transformer-based architectures, used by scBERT, scGPT, Geneformer, and scFoundation, have been successfully applied to single-cell data, leveraging the power of deep learning to model complex gene expression patterns. These models have shown promising results in various tasks, including cell-type annotation, batch integration, and multiomics analysis. However, transformervaswani2017attention has inherent limitations. The most significant one is their quadratic computational complexity with respect to sequence length, which makes them less feasible for long sequences typical of scRNA-seq data. This issue arises because the self-attention mechanism used in transformer requires computing attention scores for all pairs of tokens, leading to inefficiencies in handling long sequences. Additionally, transformer often struggles with capturing long-range dependencies in sequences, which is crucial for understanding gene regulatory networks and cell state transitions. Mambagu2023mamba has been introduced to address the transformer efficiency problem. Bidirectional Mamba(Bi-Mamba) has recently been used to solve long-sequence problems in tasks such as feature extractionsun2024bidirectional, sequential recommendationliu2024bidirectional, and time series forecastingliang2024bi. Bi-Mamba is designed to efficiently process ultra-long sequences with linear computational complexity, offering a significant reduction in memory and computation requirements compared to transformer. By leveraging state-space dynamics, Bi-Mamba can capture long-range dependencies effectively, making it well-suited for modeling gene regulatory interactions and cell state transitions. Furthermore, Bi-Mamba’s bidirectional processing enables the simultaneous consideration of upstream and downstream contexts, enhancing its ability to model complex dependencies in single-cell data. This architecture represents a promising alternative to transformer-based methods, addressing their key limitations while maintaining high performance in various single-cell analysis tasks.
@@ 3.3 Pretraining Objective
To capture the gene-gene relationships within biological pathways, we employ an InfoNCE loss function that enforces similarity among genes sharing a common pathway oord2018representation; sharma2023regenne; gundogdu2022integrating. Pathways represent functional groupings of genes that collectively contribute to specific biological processes zien2000analysis; garcia2015pathway. Genes within the same pathway are often functionally related and co-regulated, making their representations in the embedding space naturally similar. Conversely, genes that belong to different pathways should be distinct in their embeddings. To achieve this, we define gene pairs within the same pathway as positive pairs (label $y_{ij}=1$) and gene pairs from different pathways as negative pairs (label $y_{ij}=0$).
@@ 4.3 Gene Rank Reconstruction
The GeneMamba model exhibits strong capability in reconstructing ranked gene orders, a crucial requirement for single-cell analysis tasks where gene expression patterns reflect cellular states and transitions. This challenge has been highlighted in prior studies kedzierska2023assessing; boiarsky2023deep, and here we demonstrate the effectiveness of GeneMamba in addressing it. The analysis begins with the random selection of sample cells, from which input gene tokens are extracted. GeneMamba then generates output tokens, which are ranked according to their predicted likelihoods. To maintain sequence consistency, missing tokens are imputed with zeros. This strategy preserves the structural integrity of the data and enables a direct comparison between the input and output rankings.
@@ B.2 Downstream Tasks
In this experiment, we evaluate the GeneMamba’s performance to distinguish between positive pairs and negative pairs, the reference data is collected in the Gene2Vec paper [du2019gene2vec]. And we generate the embeddings using different models for positive gene pairs and negative gene pairs, then we measure the embeddings similarity of each pair. We evaluate how the distribution of positive pairs and negative pairs differ by calculating the eucilean distance, KL-Divergence, and JS-Divergence of the two distributions, the result is shown in Figure 7.
@@ B.3 Evaluation Metrics
  • GeneFormer [theodoris2023transfer]: A transformer-based model specifically designed for single-cell analysis. It uses a gene-token representation to capture complex gene-gene relationships and downstream cellular heterogeneity.
  • scGPT [cui2024scgpt]: A generative pre-trained transformer adapted for single-cell data analysis. scGPT utilizes a pretraining-finetuning paradigm to handle various single-cell tasks, including cell-type classification and clustering.
  • scFoundation [hao2024large]: A foundational model for single-cell data analysis that incorporates self-supervised learning techniques to generate robust embeddings for single-cell profiles. It excels in tasks requiring integration across batches and datasets.
  • scBERT [yang2022scbert]: A BERT-like architecture adapted for single-cell data. scBERT leverages bidirectional contextual embeddings to model complex interactions between genes within cells, making it suitable for both classification and clustering tasks.
  • Harmony [korsunsky2019fast]: A batch-effect correction method that aligns single-cell datasets from different conditions or experiments. This method is often used as a preprocessing step to ensure integration and compatibility across datasets before applying downstream machine learning methods.
@@ Appendix D Perturbation Analysis
To more comprehensively evaluate the generalizability and biological relevance of GeneMamba, we conducted a perturbation prediction study using protocols established by scGPT [cui2024scgpt]. Perturbation studies are critical for assessing a model’s ability to simulate cellular responses to genetic interventions—a fundamental requirement in functional genomics and drug discovery. We employed the Perturb-seq dataset from Norman [norman2019exploring], which include two-gene combinatorial perturbations. Specifically, we compared Mamba-based GEARS embeddings with the original GEARS model (baseline), following the benchmarking criteria previously proposed.
@@ E.1 Discretization of Gene Expression
Fixed-Size Binning: As seen in scBERT, fixed-size binning divides the range of expression values into equally sized intervals. Genes falling within the same interval are represented identically. This method, however, often leads to significant information loss, as most genes with low expression levels cluster within the same bin (e.g., the [0,1) bin [boiarsky2023deep]), masking subtle but biologically relevant differences.
@@
Value Projection Strategy: The value projection strategy, first introduced by TOSICA [chen2023transformer], has since been adopted by scCLIP [xiong2023scclip], SpaFormer [wen2023single], CellLM [zhao2023large], and CellPLM [wen2023cellplm]. In this approach, continuous gene expression values are directly projected into high-dimensional embeddings through linear or non-linear transformations. This method retains the full resolution of the data, avoiding the information loss associated with binning. However, feeding continuous values into transformer-based models diverges from traditional tokenization approaches in NLP, potentially impacting model performance. Additionally, continuous embeddings may be more sensitive to noise and batch effects, necessitating robust preprocessing steps.
Rank-Based Strategy: The rank-based strategy was first introduced in iSEEEK [shen2022universal] for single-cell transformers. This approach ranks gene expression values within each cell and retains only the top-ranked genes. The remaining genes are either truncated or represented as a lower-priority group. By focusing on the top-ranked genes, this strategy reduces sequence length, significantly improving computational efficiency. It also aligns well with the biological prioritization of highly expressed genes in specific cell states or conditions.
@@ E.2 Model Architectures for Single-Cell Analysis
Recent advancements in model architectures have addressed the computational challenges of transformer-based models in single-cell RNA sequencing (scRNA-seq) analysis, with notable innovations including scTCA [scTCA], a hybrid Transformer-CNN architecture designed for imputation and denoising of single-cell read counts to enhance downstream analyses; scHyena [scHyena], a foundation model for full-length scRNA-seq analysis in brain tissues that uses a novel transformer architecture to process data without information loss, improving tasks like cell type classification and data imputation; xTrimoGene [gong2024xtrimogene], an efficient and scalable representation learner employing an asymmetric encoder-decoder transformer architecture to reduce computational requirements while maintaining high accuracy in tasks such as cell type annotation and drug combination prediction; scTransSort [scTransSort], a transformer-based model for intelligent cell type annotation that leverages self-attention mechanisms to extract features from scRNA-seq data, achieving high accuracy and performance; scGraphformer [scGraphformer], which employs transformer-based graph neural networks to provide accurate and scalable cell type annotations, uncovering cellular heterogeneity and interactions in scRNA-seq data; scGFT [scGFT], a train-free, cell-centric generative model adept at synthesizing scRNA-seq data to enhance data augmentation and analysis; White-Box Diffusion Transformer[WhiteBoxDiffusionTransformer], a hybrid model combining diffusion models and white-box transformers to generate synthetic and biologically plausible scRNA-seq data, improving training efficiency and resource utilization; STGRNS [STGRNS], an interpretable transformer-based method for inferring gene regulatory networks from scRNA-seq data to enhance understanding of gene interactions; scRDiT [scRDiT], a generative approach using diffusion transformers to create virtual scRNA-seq data, facilitating the study of gene expression dynamics; and TransformerST [TransformerST], an unsupervised model based on the transformer architecture for super-resolution in spatial transcriptomics, improving spatial gene expression analysis. These models collectively advance the field by addressing key challenges in scRNA-seq data processing and interpretation.
@@
Traditional single-cell large language models (LLMs) are predominantly based on transformer architectures [boiarsky2023deep, white2024comprehensive]. While transformers have demonstrated exceptional performance in various single-cell transcriptomics tasks, their high computational and memory requirements make training large models time-intensive and resource-heavy. These limitations become particularly evident when processing the ultra-long sequences typical of single-cell RNA-seq data. To alleviate these challenges, some transformer-based models have incorporated optimizations to improve efficiency. For example, scBERT integrates Performer [choromanski2020rethinking] to approximate attention calculations, while scGPT employs Flash Attention [dao2022flashattention] to accelerate training by reducing memory overhead. Despite these advancements, the overall time required to train such models remains substantial, especially for datasets with millions of cells and thousands of genes. To address the intrinsic limitations of transformers, Bidirectional Mamba (Bi-Mamba) has been proposed as a novel alternative architecture. Bi-Mamba leverages state-space models (SSMs) to process ultra-long sequences with linear computational complexity in both time and memory. This represents a significant departure from traditional transformer-based methods, offering a scalable and efficient solution for the high-dimensional, high-throughput nature of single-cell transcriptomics data.

### v3

@@ 1 Introduction
Single-cell RNA sequencing (scRNA-seq) has fundamentally transformed our ability to study cellular heterogeneity and dynamics by enabling high-resolution profiling of gene expression at the individual cell level Korsunsky et al. (2019); Theodoris et al. (2023). This technology allows researchers to dissect complex cellular compositions within tissues, trace differentiation trajectories, and identify rare cell populations that remain undetectable with bulk RNA sequencing Shen et al. (2023); Wen et al. (2023b); Deconinck et al. (2021). The rich transcriptomic insights provided by scRNA-seq have catalyzed significant advancements in developmental biology, disease modeling, and drug discovery Gu and Dao (2023); Yang et al. (2024); Li and Wang (2024).
@@
Despite these advantages, the computational analysis of scRNA-seq data presents formidable challenges due to its high dimensionality, inherent sparsity, and technical noise Du et al. (2019); Zhao et al. (2023). Addressing these challenges requires robust computational models capable of capturing biologically meaningful patterns while efficiently handling the scale and variability of single-cell data. Transformer-based architectures, such as scBERT Yang et al. (2022) and scGPT Cui et al. (2024a), have emerged as powerful tools in this domain, demonstrating strong performance in tasks such as cell type classification, gene expression imputation, and differential expression analysis Bian et al. (2024); Wen et al. (2023a); Szałata et al. (2024).
@@
However, transformers exhibit critical limitations when applied to scRNA-seq data. The quadratic complexity of their self-attention mechanism constrains their scalability for long sequences, which are characteristic of single-cell transcriptomes Vaswani et al. (2017); Choromanski et al. (2020). Moreover, transformers often struggle to effectively capture long-range dependencies, which are essential for modeling gene regulatory interactions and cell state transitions Dao et al. (2022); Chen et al. (2023). These limitations have driven the exploration of alternative architectures, among which state space models (SSMs) have emerged as promising solutions, offering improved efficiency and scalability for processing long sequences Gu and Dao (2023); Dao and Gu (2024).
@@
In this study, we introduce GeneMamba, a novel state space model designed to efficiently train large-scale cell models on scRNA-seq data. SSM-based models have demonstrated competitive or superior performance compared to transformers while significantly reducing computational overhead Shen et al. (2022); Gong et al. (2024); Ryu et al. (2023). Building upon this foundation, GeneMamba incorporates bidirectional computation, enhancing its ability to capture both upstream and downstream contextual dependencies, thereby improving its performance in tasks requiring global contextual awareness Liu et al. (2024); Liang et al. (2024). We validate GeneMamba across a diverse set of applications, including multi-batch integration, cell type annotation, and gene-gene correlation analysis. Our experimental results demonstrate substantial improvements in computational efficiency and predictive performance compared to existing models Hao et al. (2024); Sun et al. (2024); Petegrosso et al. (2020). Furthermore, to assess its scalability, we conduct a comprehensive performance analysis on various variations of GeneMamba (Appendix F), highlighting its advantages in real-world biological applications.
@@ 2.1 Discretization of Gene Expression
Bin-based discretization, as used by scBERT Yang et al. (2022), scGPT Cui et al. (2024a), and scMulan Bian et al. (2024), groups expression values into predefined bins. This approach preserves absolute value distributions and simplifies sequence modeling, but may introduce information loss, particularly for genes with subtle but biologically significant expression differences. Additionally, binning can be sensitive to parameter selection, affecting downstream results.
Value projectionSzałata et al. (2024), adopted by scFoundation Hao et al. (2024) and its backbone model xTrimoGene Gong et al. (2024), projects gene expression values into continuous embeddings rather than discrete categories. This method maintains full data resolution by applying a linear transformation to the gene expression vector, which is then combined with gene-specific embeddings. However, the use of continuous embeddings diverges from traditional tokenization strategies in NLP-based transformers, and its impact on model performance remains an open question.
Rank-based discretization, utilized by Geneformer Theodoris et al. (2023), GeneCompass Yang et al. (2024), tGPTShen et al. (2023) and LangCell Zhao et al. (2024b), transforms gene expression values into ordinal rankings. This approach effectively captures relative expression levels and is more robust to batch effects and noise. Our method is based on rank-based discretization, as employed in Geneformer, which aligns naturally with biological processes such as regulatory interactions.
@@ 2.2 Model Architectures for Single-Cell Analysis
Transformer-based architectures, used by scBERT, scGPT, Geneformer, and scFoundation, have been successfully applied to single-cell data, leveraging the power of deep learning to model complex gene expression patterns. These models have shown promising results in various tasks, including cell-type annotation, batch integration, and multiomics analysis. However, transformerVaswani et al. (2017) has inherent limitations. The most significant one is their quadratic computational complexity with respect to sequence length, which makes them less feasible for long sequences typical of scRNA-seq data. This issue arises because the self-attention mechanism used in transformer requires computing attention scores for all pairs of tokens, leading to inefficiencies in handling long sequences. Additionally, transformer often struggles with capturing long-range dependencies in sequences, which is crucial for understanding gene regulatory networks and cell state transitions. MambaGu and Dao (2023) has been introduced to address the transformer efficiency problem. Bidirectional Mamba(Bi-Mamba) has recently been used to solve long-sequence problems in tasks such as feature extractionSun et al. (2024), sequential recommendationLiu et al. (2024), and time series forecastingLiang et al. (2024). Bi-Mamba is designed to efficiently process ultra-long sequences with linear computational complexity, offering a significant reduction in memory and computation requirements compared to transformer. By leveraging state-space dynamics, Bi-Mamba can capture long-range dependencies effectively, making it well-suited for modeling gene regulatory interactions and cell state transitions. Furthermore, Bi-Mamba’s bidirectional processing enables the simultaneous consideration of upstream and downstream contexts, enhancing its ability to model complex dependencies in single-cell data. This architecture represents a promising alternative to transformer-based methods, addressing their key limitations while maintaining high performance in various single-cell analysis tasks.
@@ 3.3 Pretraining Objective
To capture the gene-gene relationships within biological pathways, we employ an InfoNCE loss function that enforces similarity among genes sharing a common pathway Oord et al. (2018); Sharma and Xu (2023); Gundogdu et al. (2022). Pathways represent functional groupings of genes that collectively contribute to specific biological processes Zien et al. (2000); García-Campos et al. (2015). Genes within the same pathway are often functionally related and co-regulated, making their representations in the embedding space naturally similar. Conversely, genes that belong to different pathways should be distinct in their embeddings. To achieve this, we define gene pairs within the same pathway as positive pairs (label $y_{ij}=1$) and gene pairs from different pathways as negative pairs (label $y_{ij}=0$).
@@ 4.3 Gene Rank Reconstruction
The GeneMamba model exhibits strong capability in reconstructing ranked gene orders, a crucial requirement for single-cell analysis tasks where gene expression patterns reflect cellular states and transitions. This challenge has been highlighted in prior studies Kedzierska et al. (2023); Boiarsky et al. (2023), and here we demonstrate the effectiveness of GeneMamba in addressing it. The analysis begins with the random selection of sample cells, from which input gene tokens are extracted. GeneMamba then generates output tokens, which are ranked according to their predicted likelihoods. To maintain sequence consistency, missing tokens are imputed with zeros. This strategy preserves the structural integrity of the data and enables a direct comparison between the input and output rankings.
@@ B.2 Downstream Tasks
In this experiment, we evaluate the GeneMamba’s performance to distinguish between positive pairs and negative pairs, the reference data is collected in the Gene2Vec paper [11]. And we generate the embeddings using different models for positive gene pairs and negative gene pairs, then we measure the embeddings similarity of each pair. We evaluate how the distribution of positive pairs and negative pairs differ by calculating the eucilean distance, KL-Divergence, and JS-Divergence of the two distributions, the result is shown in Figure 7.
@@ B.3 Evaluation Metrics
  • GeneFormer [35]: A transformer-based model specifically designed for single-cell analysis. It uses a gene-token representation to capture complex gene-gene relationships and downstream cellular heterogeneity.
  • scGPT [5]: A generative pre-trained transformer adapted for single-cell data analysis. scGPT utilizes a pretraining-finetuning paradigm to handle various single-cell tasks, including cell-type classification and clustering.
  • scFoundation [17]: A foundational model for single-cell data analysis that incorporates self-supervised learning techniques to generate robust embeddings for single-cell profiles. It excels in tasks requiring integration across batches and datasets.
  • scBERT [42]: A BERT-like architecture adapted for single-cell data. scBERT leverages bidirectional contextual embeddings to model complex interactions between genes within cells, making it suitable for both classification and clustering tasks.
  • Harmony [20]: A batch-effect correction method that aligns single-cell datasets from different conditions or experiments. This method is often used as a preprocessing step to ensure integration and compatibility across datasets before applying downstream machine learning methods.
@@ Appendix D Perturbation Analysis
To more comprehensively evaluate the generalizability and biological relevance of GeneMamba, we conducted a perturbation prediction study using protocols established by scGPT [5]. Perturbation studies are critical for assessing a model’s ability to simulate cellular responses to genetic interventions—a fundamental requirement in functional genomics and drug discovery. We employed the Perturb-seq dataset from Norman [24], which include two-gene combinatorial perturbations. Specifically, we compared Mamba-based GEARS embeddings with the original GEARS model (baseline), following the benchmarking criteria previously proposed.
@@ E.1 Discretization of Gene Expression
Fixed-Size Binning: As seen in scBERT, fixed-size binning divides the range of expression values into equally sized intervals. Genes falling within the same interval are represented identically. This method, however, often leads to significant information loss, as most genes with low expression levels cluster within the same bin (e.g., the [0,1) bin [2]), masking subtle but biologically relevant differences.
@@
Value Projection Strategy: The value projection strategy, first introduced by TOSICA [3], has since been adopted by scCLIP [40], SpaFormer [38], CellLM [46], and CellPLM [37]. In this approach, continuous gene expression values are directly projected into high-dimensional embeddings through linear or non-linear transformations. This method retains the full resolution of the data, avoiding the information loss associated with binning. However, feeding continuous values into transformer-based models diverges from traditional tokenization approaches in NLP, potentially impacting model performance. Additionally, continuous embeddings may be more sensitive to noise and batch effects, necessitating robust preprocessing steps.
Rank-Based Strategy: The rank-based strategy was first introduced in iSEEEK [32] for single-cell transformers. This approach ranks gene expression values within each cell and retains only the top-ranked genes. The remaining genes are either truncated or represented as a lower-priority group. By focusing on the top-ranked genes, this strategy reduces sequence length, significantly improving computational efficiency. It also aligns well with the biological prioritization of highly expressed genes in specific cell states or conditions.
@@ E.2 Model Architectures for Single-Cell Analysis
Recent advancements in model architectures have addressed the computational challenges of transformer-based models in single-cell RNA sequencing (scRNA-seq) analysis, with notable innovations including scTCA [44], a hybrid Transformer-CNN architecture designed for imputation and denoising of single-cell read counts to enhance downstream analyses; scHyena [26], a foundation model for full-length scRNA-seq analysis in brain tissues that uses a novel transformer architecture to process data without information loss, improving tasks like cell type classification and data imputation; xTrimoGene [14], an efficient and scalable representation learner employing an asymmetric encoder-decoder transformer architecture to reduce computational requirements while maintaining high accuracy in tasks such as cell type annotation and drug combination prediction; scTransSort [18], a transformer-based model for intelligent cell type annotation that leverages self-attention mechanisms to extract features from scRNA-seq data, achieving high accuracy and performance; scGraphformer [12], which employs transformer-based graph neural networks to provide accurate and scalable cell type annotations, uncovering cellular heterogeneity and interactions in scRNA-seq data; scGFT [25], a train-free, cell-centric generative model adept at synthesizing scRNA-seq data to enhance data augmentation and analysis; White-Box Diffusion Transformer[6], a hybrid model combining diffusion models and white-box transformers to generate synthetic and biologically plausible scRNA-seq data, improving training efficiency and resource utilization; STGRNS [41], an interpretable transformer-based method for inferring gene regulatory networks from scRNA-seq data to enhance understanding of gene interactions; scRDiT [10], a generative approach using diffusion transformers to create virtual scRNA-seq data, facilitating the study of gene expression dynamics; and TransformerST [45], an unsupervised model based on the transformer architecture for super-resolution in spatial transcriptomics, improving spatial gene expression analysis. These models collectively advance the field by addressing key challenges in scRNA-seq data processing and interpretation.
@@
Traditional single-cell large language models (LLMs) are predominantly based on transformer architectures [2, 39]. While transformers have demonstrated exceptional performance in various single-cell transcriptomics tasks, their high computational and memory requirements make training large models time-intensive and resource-heavy. These limitations become particularly evident when processing the ultra-long sequences typical of single-cell RNA-seq data. To alleviate these challenges, some transformer-based models have incorporated optimizations to improve efficiency. For example, scBERT integrates Performer [4] to approximate attention calculations, while scGPT employs Flash Attention [7] to accelerate training by reducing memory overhead. Despite these advancements, the overall time required to train such models remains substantial, especially for datasets with millions of cells and thousands of genes. To address the intrinsic limitations of transformers, Bidirectional Mamba (Bi-Mamba) has been proposed as a novel alternative architecture. Bi-Mamba leverages state-space models (SSMs) to process ultra-long sequences with linear computational complexity in both time and memory. This represents a significant departure from traditional transformer-based methods, offering a scalable and efficient solution for the high-dimensional, high-throughput nature of single-cell transcriptomics data.